Given: mean is $25 and standard deviation is $8. Find dollar amount such that 80% of all customers spend atleast this amount.
Answer by stanbon(75887)

Given: mean is $25 and standard deviation is $8. Find dollar amount such that 80% of all customers spend at least this amount.
---
Find the z-value with a left tail of 20%.
invNorm(0.2) = -0.8416
---
Find the corresponding x-value:
x = zs + u
x = -0.8416*8 + 25 = $18.27
=================
Cheers,
stan H.